what is an angle's measure is 30 degrees more than its supplement.
Answer #1
X+(X+30)=180 2X=180-30 2X=150 X=150/2 X=75 ANSWER FOR THE SMALLER SUPPLEMENTAL ANGLE. 75+30=105 FOR THE LARGER ANGLE. PROOF 75+105=180 180=180
answered by: diana
Answer #2
what is an angle's measure is 30 degrees more than its supplement. ------------------------- Let the angle be "x". The supplement = 180-x ----------------------- EQUATION: x = 180-x+30 2x = 210 x = 105 degrees ---------------- Cheers, Stan H.
answered by: umair